Uri Onn: Representations of GL(n, O) with geometric origin

Let O be the ring of integers of a non-archimedean local field. For example, O can be a ring of power series 𝔽_{p}[[x]] or the ring of p-adic integers ℤ_{p}. We shall discuss complex continuous representations of the group GL(n, O). The main emphasis is on representations arising in a geometric context, that is, from the action of the group on lattices of submodules.